OpenStudy (anonymous):

use the method of substitution to solve. 4a=2-b 4a=4-b

OpenStudy (anonymous):

i believe that there is not a solution but want to be sure

jimthompson5910 (jim_thompson5910):

4a=2-b 4 - b=2-b 4 = 2 so you are correct

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

yw